# Homogenized C. elegans Neural Activity and Connectivity Data

### v2

@@ Abstract
There is renewed interest in modeling and understanding the nervous system of the nematode Caenorhabditis elegans (C. elegans). This is particularly interesting as this model system provides a path to bridge the gap between structure and function, from nervous system connectivity to physiology. However, the many existing physiology datasets, both recording and stimulation, as well as connectome datasets, are in distinct formats, requiring extra processing steps before modeling or other analysis can commence. Here we present a homogenized dataset of neural activity, including during stimulation, compiled from 11 neuroimaging experiments and from 10 connectome reconstructions. The physiology datasets, collected under varying experimental protocols, all measure neural activity via calcium fluorescence in labeled subsets of the worm’s 300 neurons. Our preprocessing pipeline standardizes these datasets by consistently ordering labeled neurons and resampling traces to a common sampling frequency. The resulting dataset includes neural recordings from approximately 900 worms and 250 uniquely labeled neurons. The connectome datasets, collected from electron microscopy (EM) reconstructions, all contain the entire nervous system of the worm, preprocessed into a graph of connections across the neurons. Using our collection of datasets is facilitated through easy data sharing on HuggingFace. We believe that our joint dataset of physiology and connectivity will facilitate modeling, for example in terms of recurrent neural network or transformer architectures, making it easier to check how well models generalize across animals and labs.
@@ 1 Introduction
Understanding neural dynamics and their relationship to underlying structure remains one of the key challenges in neuroscience and machine learning. While neural dynamics in the brain can often be complex and difficult to measure in larger organisms, smaller systems like C. elegans offer a unique opportunity to create comprehensive models that bridge functional activity with known connectomic structure. C. elegans has a relatively small nervous system of 300 neurons [17], and both its neural dynamics and complete synaptic connectome have been extensively studied and mapped out [24, 3, 25]. This makes it an ideal organism for studying how neural dynamics arise from connectivity and how this relationship could inform both neuroscience and AI models of neural function.
With its transparent body and well-characterized neural architecture, C. elegans provides an excellent model for whole-brain functional imaging using calcium fluorescence sensors [26]. Calcium indicators act as proxies for local computation, allowing researchers to capture neural activity at single-neuron resolution across many neurons simultaneously [6]. Additionally, since the full connectome of C. elegans has been mapped via electron microscopy [24, 3, 25], combining these two datasets — neural dynamics and connectivity structure — allows for a deep investigation of how network connectivity influences neural function. Unification of such datasets can also contribute to the broader goal of understanding how neural activity propagates through networks, ultimately informing both the development of AI-based neural systems and providing insights into how these mechanisms generalize to more complex organisms [8].
However, the existing datasets from calcium imaging studies of C. elegans are often recorded under various experimental conditions, with different subsets of neurons labeled and differing sampling rates. This makes it challenging to perform cross-study comparisons or build unified models.
To address these challenges, we have created two datasets, the first combining neural dynamics from 11 calcium maging studies and the second combining neural connectivity from 3 connectome studies of C. elegans [24, 3, 25]. This dataset integrates both structure and function, making it an invaluable resource for computational modelers interested in building neural network models grounded in real biological data. The primary goal of this dataset is to facilitate the training of computational models of C. elegans neural dynamics, which can be used to study how neural function is related to structure. The dataset includes preprocessed neural activity data that has been normalized, resampled to a common time step, and smoothed using a causal exponential kernel. It also includes graph-based representations of the connectome, with detailed information about neuron positions and synaptic connections, facilitating the development of models that combine structure and function. We are releasing these datasets as open-source resources on the HuggingFace platform to ensure broad accessibility to researchers in both neuroscience and machine learning. By unifying these datasets and making them accessible, we hope to encourage the development of computational models that can capture the structure-function relationships in small nervous systems.
@@ 2.1 Code Repository
We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that implement the entire preprocessing pipeline we describe in following sections. Our preprocessing is combination of utilities (classes, functions and files) that facilitate the extraction and processing of a collection of C. elegans neural and connectome data in a standardized format.
The main classes in the repository are NeuralBasePreprocessor and ConnectomeBasePreprocessor. The former manages processing of neural activity data as measured by calcium fluorescence imaging [23]. The latter manages the processing of synaptic connectivity data as measured by anatomical counts from electron microscopy [24].
The main functions in the repository are pickle_neural_data() and preprocess_connectome(). The former handles the extraction and processing of the raw calcium data from each of our 11 neural activity source datasets into standardized neural activity data that are saved as compressed files. The latter processes the raw connectome data from connectome source files into compressed graph tensors that are saved as compressed in a data format amenable to downstream processing by graph neural network packages. There are 10 distinct connectome source files as shown in Table 3 but they are all derived from 4 primary source publications [24, 3, 25, 14].
@@ 2.2 Calcium Fluorescence Data
We processed neural activity datasets from 11 open-source studies of C. elegans, each measuring calcium fluorescence ($\Delta F/F_{0}$) in subsets of the worm’s neurons under various conditions [14, 5, 18, 26, 9, 15, 12, 10, 11, 13]. These datasets include a variety of experimental protocols, such as freely moving, immobilized, and optogenetically stimulated animals. The number of worms and identified neurons varies across datasets as showb in Table 1). All animals were hermaphrodites at developmental stages from L4 to adulthood.
Table 1: Calcium Fluorescence Neural Activity Datasets Metadata. Metadata for calcium fluorescence neural activity datasets collected from various sources. Each dataset includes information about the experimental conditions, the number of worms, the average number of labeled and recorded neurons, and the range of labeled neurons observed.
Source Dataset Name Database Link Data Files Num. Worms Mean Num. Neurons (labeled, recorded) Num. labeled neurons (min, max)
Kato2015 [9] osf.io/2395t/ WT_Stim.mat 12 (42, 127) (31, 51)
Nichols2017 [12] osf.io/kbf38/ let.mat 44 (34, 108) (23, 43)
Skora2018 [15] osf.io/za3gt/ WT_.mat 12 (46, 129) (39, 55)
Kaplan2020 [9] osf.io/9nfhz/ Neuron2019_Data_ *.mat 19 (36, 114) (23, 51)
Yemini2021 [26] zenodo.org/records/ 3906530 _Activity_OH.mat 49 (110, 125) (33, 179)
Uzel2022 [18] osf.io/3vkxn Uzel_WT.mat 6 (50, 138) (46, 58)
Dag2023 [13] tinyurl.com/github Dag2023 data/swf702_with_ id/*.h5 7 (100, 143) (87, 110)
Atanas2023 [5] wormwideweb.org YYYY-MM-DD-.json/h5 42 (88, 136) (64, 115)
Leifer2023 [14] tinyurl.com/drive Leifer2023 exported_data.tar.gz 108 (64, 68) (19, 98)
Lin2023 [11] tinyurl.comdropbox Lin2023 run_prfrd_data.mat 577 (8, 8) (1, 22)
Venkatachalam2024 [22] chemosensory-data.worm.world 2022_herm_.zip 22 (187, 187) (185, 189)
@@ 2.2.1 Neural Data Processing
Table 2: Notation Used in the Neural Data Processing Pipeline. Description of the mathematical notation used throughout the neural data preprocessing pipeline. Variables represent worm-specific indices, neural activity data structures, temporal properties, neuron counts, and preprocessing parameters.
Symbol Description
$k$ Worm index variable within a single dataset
$\mathbf{X}^{(k)}$ Neural activity data for a single worm $k$
$T_{k}$ Number of time points of data recorded for a worm $k$
$D$ Number of neurons in C. elegans
$\mathbf{\mu},\mathbf{\sigma}$ Mean and standard deviation over time of neural activity data for each neuron
$\alpha$ Optional smoothing hyperparameter
$\mathbf{M}^{(k)}$ Binary mask generated for each worm using labeled neurons applied to neural activity data
@@
2. Custom Dataset Implementations. Each source dataset has a custom class that inherits from the NeuralBasePreprocessor parent class, facilitating data extraction for each unique format. The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps: normalization, smoothing, resampling, and masking.
@@
3. Neuron Trace Normalization. The first preprocessing step is to normalize or z-score the raw (i.e. from the source datasets) neural activity across the neuron dimension. For each neuron trace $i$ in worm $k$, we compute:
@@
where $\mathbf{\mu},\mathbf{\sigma}\in\mathbb{R}^{D}$ are the mean and standard deviation over time for each neuron.
@@
6. Masking and Subsetting. We also create a binary mask for each worm, $\mathbf{M}^{(k)}\in\{0,1\}^{D}$, where $D=300$ represents the total number of canonical neurons in C. elegans. Each element $\mathbf{M}^{(k)}_{i}=1$ indicates that neuron $i$ was labeled and measured for worm $k$, while $\mathbf{M}^{(k)}_{i}=0$ indicates that neuron $i$ was not measured.
The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$ enabling in its potential use in downstream analyses that require filtering out unlabeled neurons. Therefore, the pair $(\mathbf{X}^{(k)},\mathbf{M}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
@@
Figure 1: Processed Neural Activity Data Structure. Illustration of the structure of our neural activity data as a time-series matrix, where each column represents a neuron, ordered alphabetically by their canonical names in C. elegans. The calcium signals are normalized, smoothed, and resampled to a common time interval ($\Delta t$). A binary mask is generated to indicate which neurons were labeled for each worm. The mask does not modify the data but is retained as metadata to aid future analyses (e.g., selecting only labeled neurons).
@@
         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\mathbf{X}^{(k)}_{j}\mid\mathbf{X}^{(k)}_{j}\neq\emptyset,\,j\in[1,D]\}$
@@
Combine the processed datasets
@@ 2.3 Connectome Graph Data
We processed multiple open-source C. elegans connectome datasets that capture the full neural wiring diagram, including both chemical synapses and gap junctions. Table 5 summarizes these datasets, which vary in annotation precision and recording techniques.
Table 3: Open-Source Connectome Datasets Metadata.Metadata for open-source C. elegans connectome datasets. Each dataset was standardized into the graph_tensor format, listing the corresponding graph tensor file and the total number of edges (synaptic and electrical connections) per dataset.
Source Dataset Name Graph Tensor File Number of Edges
Cook2019 [3] graph_tensors_cook2019.pt 46837
White1986 (whole) [24] graph_tensors_white1986_whole.pt 46603
White1986 (n2u) [24] graph_tensors_white1986_n2u.pt 45808
White1986 (jsh) [24] graph_tensors_white1986_jsh.pt 45759
White1986 (jse) [24] graph_tensors_white1986_jse.pt 44901
Witvliet2020 (7) [25] graph_tensors_witvliet2020_7.pt 46057
Witvliet2020 (8) [25] graph_tensors_witvliet2020_8.pt 46107
OpenWorm2023 [21] graph_tensors_openworm.pt 47683
Randi2023 [14] graph_tensors_funconn.pt 50521
Chklovskii2011 [19] graph_tensors_chklovskii.pt 46655
The connectome data is represented as a directed graph where neurons are nodes and edges represent synaptic or gap junction connections.
Figure 2: Graphical Representation of the Connectome Data. Graph-based structure used for the C. elegans connectome data. Neurons are represented as nodes, with connections (edges) classified as either chemical synapses or gap junctions. The neuron features, such as positional coordinates, neuron class, and type, are stored as node attributes, while edge attributes capture synaptic weights and connection types. This graph-based structure enables connectivity-informed modeling of neural activity. The differently colored nodes arbitrarily represent the presence of different types of neurons within the C. elegans nervous system (e.g. motor, inter & sensory neurons). The tables correspond to the bolded edge and node. The attr(e_1) feature for a single edge is the tuple of chemical synapse and gap junction weights.
2.3.1 Graph Tensor Data Processing

### v3

@@ Abstract
There is renewed interest in modeling and understanding the nervous system of the nematode Caenorhabditis elegans (C. elegans), as this small model system provides a path to bridge the gap between nervous system structure (connectivity) and function (physiology). However, existing physiology datasets, whether involving passive recording or stimulation, are in distinct formats, and connectome datasets require preprocessing before analysis can commence. Here we compile and homogenize datasets of neural activity and connectivity. Our neural activity dataset is derived from 11 C. elegans neuroimaging experiments, while our connectivity dataset is compiled from 9 connectome annotations based on 3 primary electron microscopy studies and 1 signal propagation study. Physiology datasets, collected under varying protocols, measure calcium fluorescence in labeled subsets of the worm’s 300 neurons. Our preprocessing pipeline standardizes these datasets by consistently ordering labeled neurons and resampling traces to a common sampling rate, yielding recordings from approximately 900 worms and 250 uniquely labeled neurons. The connectome datasets, collected from electron microscopy reconstructions, represent the entire nervous system as a graph of connections. Our collection is accessible on HuggingFace, facilitating analysis of the structure-function relationship in biology using modern neural network architectures and enabling cross-lab and cross-animal comparisons.
@@ 1 Introduction
Understanding neural dynamics and their relationship to underlying structure remains one of the key challenges in neuroscience. Recent advances in computational modeling and data acquisition have reinvigorated interest in the neural system of the nematode Caenorhabditis elegans (C. elegans), a model organism uniquely suited for bridging the gap between neural connectivity and function. With its fully mapped synaptic connectome [35, 36], consisting of approximately 300 neurons [3], and its extensively studied neuronal dynamics [23, 37], C. elegans provides an ideal platform for investigating how neural structure constrains or influences physiology.
Recent research has focused on leveraging the C. elegans connectome for generative models of neural development [30] and linking whole-brain activity to synaptic connectivity [4]. These studies exemplify its potential for elucidating how neural signaling is determined or constrained by connectivity and for inspiring biologically grounded computational frameworks [26]. Furthermore, the transparent body of C. elegans, combined with its well-characterized neural architecture, makes it an excellent candidate for whole-brain functional imaging using calcium fluorescence sensors [37]. Calcium imaging enables the capture of neural dynamics at single-neuron resolution across many neurons simultaneously [6], serving as a proxy for neuronal activity. Although calcium fluctuations may primarily reflect local processes in neurites rather than whole-neuron computations [24], compact neurons like those in C. elegans likely rely on such localized signaling mechanisms, making this an informative approach.
The combination of calcium imaging datasets with the full connectome, mapped via electron microscopy [35, 3, 36], offers an unprecedented opportunity to investigate how connectivity influences neural function. Unified datasets of structure and function can provide valuable insights into neural activity propagation, informing both the development of artificial intelligence (AI)-based neural simulations and the understanding of mechanisms that generalize to more complex organisms [8].
However, calcium imaging datasets from C. elegans experiments vary widely in their experimental conditions, labeled neurons, and sampling rates, complicating cross-study comparisons and unified modeling efforts. Similarly, connectome datasets—whether physical or functional—require significant preprocessing to facilitate analysis or computational modeling.
To address these challenges, we have created two integrated datasets. The first aggregates neural dynamics from 11 calcium imaging studies, and the second compiles connectome data from three electron microscopy wiring studies [35, 3, 36] and one functional signal propagation study [14]. Together, these datasets combine structure and function, offering an invaluable resource for computational modelers aiming to construct biologically grounded neural network models.
Our preprocessing pipeline standardizes neural activity data through normalization, resampling to a common rate, and consistent organization into data structures. Neural recordings include data from approximately 900 worms and 250 uniquely labeled neurons. The connectome data is represented as a graph with attributes for neuron positions, synaptic connections, and additional metadata, facilitating downstream analyses that integrate structure and function.
These datasets are available as open-source resources on the HuggingFace platform, ensuring accessibility for researchers across neuroscience and machine learning. By unifying and releasing these datasets, we aim to catalyze the development of computational models that capture the structure-function relationships in small neural systems, enabling robust cross-lab and cross-animal comparisons.
@@ 2.1 Code Repository
We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that implement the entire preprocessing pipeline we describe in following sections. Our preprocessing is combination of utilities (classes, functions, files and links) that facilitate the extraction and processing of a collection of C. elegans neural and connectome data into standardized formats.
The main classes in the repository are NeuralBasePreprocessor and ConnectomeBasePreprocessor. The former manages processing of neural activity data as measured by calcium fluorescence imaging [33]. The latter manages the processing of synaptic connectivity data as measured by anatomical counts from electron microscopy (EM) [35].
The main functions in the repository are pickle_neural_data() and preprocess_connectome(). The pickle_neural_data() function extracts and processes raw calcium fluorescence data from each of our 11 neural activity source datasets, standardizing them into neural activity matrices that are saved as compressed files for efficient storage and retrieval. The preprocess_connectome() function processes raw connectome data from source files into graph tensors, which are formatted for compatibility with graph neural network (GNN) frameworks such as PyTorch Geometric [21]. This preprocessing step facilitates downstream tasks like connectivity-informed modeling of neural activity. Although there are 10 distinct connectome source files, as detailed in Table 5, all are derived from 4 primary publications [35, 3, 36, 14], ensuring comprehensive coverage of C. elegans neural connectivity data.
@@
Figure 1: Overview of Neural Activity Datasets. (A) Distribution of the number of worms in each experimental dataset. (B) Number of recorded neurons per worm compared to the total neuronal population of C. elegans. (C) Total duration of recorded neural activity for each dataset. (D) Average duration of recorded activity per worm, with one hour of calcium imaging as a benchmark. (E) Number of sampled time steps per worm, and (F) Sampling intervals for recorded neural activity. The horizontal dashed lines in (E) and (F) indicate the resampled sequence length ($L=314$) and time step ($\Delta t=0.333$ seconds) used in our preprocessing pipeline.
@@ 2.2 Calcium Fluorescence Data
We processed neural activity datasets from 11 open-source studies of C. elegans, each measuring calcium fluorescence ($\Delta F/F_{0}$) in subsets of the worm’s neurons under various conditions [14, 5, 18, 37, 9, 15, 12, 10, 11, 13]. These datasets include a variety of experimental protocols, such as freely moving, immobilized, and optogenetically stimulated animals. The number of worms and identified neurons varies across datasets as shown in Table 1. All worms were hermaphrodites at developmental stages no earlier than L4 [29], with the most being in adulthood (Figure 5).
Table 1: Calcium Fluorescence Neural Activity Datasets Metadata. Metadata for calcium fluorescence neural activity datasets collected from various sources. The table includes the number of worms, the average number of labeled and recorded neurons, the range of labeled neurons observed, and summary experimental protocol details.
Index Dataset Name Num. Worms Mean Num. Neurons (labeled, recorded) Num. Labeled Neurons (min, max) Experimental Protocol
1 Kato2015 [10] 12 (42, 127) (31, 51) Wild-type; L4 stage; immobilized in a microfluidic device
2 Nichols2017 [12] 44 (34, 108) (23, 43) Wild-type and npr-1 mutants; Lethargus/pre-lethargus stages; immobilized
3 Skora2018 [15] 12 (46, 129) (39, 55) Wild-type; fasting/starvation conditions; immobilized
4 Kaplan2020 [9] 19 (36, 114) (23, 51) Wild-type; L4 stage; freely moving and immobilized
5 Yemini2021 [37] 49 (110, 125) (33, 179) Transgenic strain OH15500; adult stage; immobilized
6 Uzel2022 [18] 6 (50, 138) (46, 58) Well-fed wild-type; adult stage; immobilized
7 Dag2023 [13] 7 (100, 143) (87, 110) Wild-type; adult stage; freely moving with subsequent NeuroPAL immobilization
8 Atanas2023 [5] 42 (88, 136) (64, 115) Transgenic strain NLS-GCaMP7f; adult stage; freely moving
9 Leifer2023 [14] 103 (69, 122) (26, 102) Wild-type; adult stage; immobilized
10 Lin2023 [11] 577 (8, 8) (1, 22) Wild-type; adult stage; immobilized
11 Venkatachalam2024 [32] 22 (187, 187) (185, 189) Wild-type; adult stage; freely moving
2.2.1 Overview of Neural Activity Datasets
We analyzed neural activity datasets collected from 11 independent studies, which varied in terms of the number of worms recorded, the number of neurons labeled, the duration of recordings, and sampling intervals. Figure 1 provides an overview of these datasets, highlighting their variability in experimental conditions.
@@ 2.2.2 Neural Data Processing
2. Custom Dataset Implementations. Each source dataset has a custom class that inherits from the NeuralBasePreprocessor parent class, facilitating data extraction for each unique format. The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps: normalization, smoothing, resampling, and masking. Adding a new source dataset to the pipeline is as easy as defining a new child preprocessor class that implements any loading and processing logic that needs to be customized to that particular source; for example, if the new source’s data is stored in a special file type.
Figure 2: Processed Neural Activity Data Structure. Illustration of the structure of our neural activity data as a time-series matrix, where each column represents a neuron, ordered alphabetically by their canonical names in C. elegans. The calcium signals are normalized, (optionally) smoothed, and resampled to a common time interval ($\Delta t$). A binary mask is generated to indicate which neurons were labeled for each worm.
@@
3. Neuron Trace Normalization. The first preprocessing step is to normalize ($z$-score) the raw neural activity (i.e. data loaded from the source dataset files) across the neuron dimension. For each neuron trace $i$ in worm $k$, we compute:
@@
where $\boldsymbol{\mu},\boldsymbol{\sigma}\in\mathbb{R}^{D}$ are vectors of the mean and standard deviation, respectively, for each neuron, computed across the temporal dimension.
@@
6. Masking and Subsetting. We also create a binary mask for each worm, $\mathbf{M}^{(k)}\in\{0,1\}^{D}$, where $D=300$ represents the total number of canonical neurons in C. elegans. Each element $\mathbf{M}^{(k)}_{i}=1$ indicates that neuron $i$ was labeled in worm $k$, while $\mathbf{M}^{(k)}_{i}=0$ indicates that neuron $i$ was not labeled.
In some of the source datasets we processed there were instances of calcium fluorescence measurements from neurons which were not identified with a label. For simplicity, we will not make a distinction in this work between unlabeled neurons and unmeasured neurons and may use the two terms interchangeably.
The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$. We do not apply the mask to modify the neural data matrix (e.g. zeroing out measurements for unlabeled neurons) but instead retain it as a secondary data structure that can be used in downstream modeling and analysis (e.g., selecting only labeled neurons). Therefore, the pair $(\mathbf{X}^{(k)},\mathbf{M}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
@@
         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\texttt{boolean}(\mathbf{X}^{(k)}_{j}\neq\emptyset)\mid j\in[1,D]\}$
@@
Compile and store all processed datasets $\mathcal{D}_{\mathrm{processed,all\_sources}}$
@@ 2.3 Connectome Graph Data
We processed multiple open-source C. elegans connectome datasets, each capturing the neural wiring diagram, including both chemical synapses and electrical synapse (also referred as gap junctions). Table 5 provides metadata for these datasets, which vary in annotation precision, formats, and the extent of the nervous system covered.
2.3.1 Connectome Representation
The connectome is represented as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where:
1. Nodes ($\mathcal{V}$): Represent individual neurons.
$\mathcal{V}:=\{\nu_{i}\}_{i=1}^{D},\quad D:=|\mathcal{V}|=300.$
Each node $\nu_{i}$ is associated with attributes, such as:
  • Positional Coordinates: Spatial location of the neuron in body-atlas coordinates $(x,y,z)$ [16].
  • Neuron Class and Type: Categories such as sensory, motor, or interneuron, and neuron-specific types (e.g., ADA, IL2, VD).
  • Neurotransmitter: Identity of neurotransmitters released by the neuron (e.g., acetylcholine, dopamine) [27].
  • Feature Vector ($\mathbf{x}_{i}$): Optionally, numerical data such as calcium fluorescence measurements over time.
2. Edges ($\mathcal{E}$): Represent synaptic and electrical connections between neurons.
:=$\mathcal{E}:=\{e_{ij}\}_{k=1}^{E},\quad E:=|\mathcal{E}|.$
Each edge $e_{ij}$ is associated with an attribute vector $\mathbf{a}_{ij}$, which can include:
  • $c$: Chemical synapse weight.
  • $g$: Gap junction weight.
  • $f$: Functional connectivity strength (optional, based on optogenetic perturbations [14]).
The neighbors of a node $\nu_{i}$ are defined as $\mathcal{N}(\nu_{i}):=\{\nu_{j}:e_{ij}\vee e_{ji}\in\mathcal{E}\}$, where $\vee$ indicates "or", ignoring directionality.
Edges are included in $\mathcal{E}$ if any of their attributes $\mathbf{a}_{ij}$ are nonzero. For example, an edge may represent gap junctions ($\mathbf{a}_{ij}=[0,g]$) or chemical synapses ($\mathbf{a}_{ij}=[c,0]$). In cases where functional connectivity is included, the graph may become fully connected ($|\mathcal{E}|\approx D^{2}$) unless significance thresholds are applied. Chemical and electrical synapse weights are based on counts from EM annotations [35, 3, 36] whereas the functional connectivity is based on signal propagation [14].
Figure 3 illustrates this graph-based representation of the nervous system, highlighting node and edge attributes. Table 5 summarizes the notation used in constructing and describing this graph structure.
2.3.2 Data Extraction and Preprocessing